OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 10 = x2 + 3x? another study guide question for the asvab? (Armed Services Vocational Aptitude Battery)

OpenStudy (lalaly):

x intercepts are when y=0

OpenStudy (anonymous):

hello again lalaly

OpenStudy (lalaly):

:) hi gatorgavlet

OpenStudy (anonymous):

:)

OpenStudy (lalaly):

substitute y=0 and solve for x ...

OpenStudy (anonymous):

oh okay hold on a sec

OpenStudy (lalaly):

oky

OpenStudy (anonymous):

(-2, 0) and (5, 0)? This is what I keep getting but it doesn't seem right

OpenStudy (lalaly):

setting y=0 0+10=x^2+3x x^2+3x-10=0 (x+5)(x-2)=0 x=-5,2 the x intercepts are (-5,0) and (2,0)

OpenStudy (anonymous):

so i had the numbers rite, just not the signs

OpenStudy (lalaly):

lol yes

OpenStudy (anonymous):

thx again :)

OpenStudy (lalaly):

anytime

OpenStudy (anonymous):

i need to study more often lol my mind just doesn't like algebra

OpenStudy (lalaly):

no one does:P